# Dicke states as matrix product states

### v1

@@ Appendix A Sequential preparation of Dicke states
$I^{(i)}_{l}\,|\underline{j}\rangle\,|0\rangle_{i}=\begin{cases}|\underline{j}\rangle\,|0\rangle_{i}&l<j\\[7.22743pt] \sum_{m=0}^{1}\gamma^{(i)}_{j,m}\,|\underline{j+m}\rangle\,|m\rangle_{i}&l=j\end{cases}\,;$ (A.2)
moreover, the $I^{(i)}_{l}$ operators are ordered so that they do not interfere with each other:
@@
since $\mathcal{U}\,|\underline{0}\rangle\,|0\rangle^{\otimes n}=|\underline{k}\rangle\,|D^{n}_{k}\rangle$, see (2.4) and (3.3). It is possible to show that the product over $l$ in (A.1) can be restricted, so that $U_{i}$ is given by
@@
Let us represent $U_{i}$ (A.7) by the circuit diagram in Fig. 2(a), where the top wire represents qubit $i$, and the bottom wire represents the $(k+1)$-level qudit ancilla. The operator $\mathcal{U}$ (A.6) can then be represented by the circuit diagram in Fig. 2(b). The size and depth of the circuit is $3k(n+1-k)$, as follows from (A.6) and (A.7). Code in cirq [41] for simulating this circuit is provided in the Supplementary Material.
[…]
(a) $U_{i}=\prod_{l}I^{(i)}_{l}$, with $x=\max(0,i-n+k-1)$ and $y=\min(i-1,k-1)$
@@
(b) $\mathcal{U}=\prod_{i}U_{i}$
Figure 2: Circuit diagram for sequential Dicke state preparation

### v2

@@ Appendix A Sequential preparation of Dicke states
$I^{(i)}_{l}\,|\underline{j}\rangle\,|0\rangle_{i}=\begin{cases}|\underline{j}\rangle\,|0\rangle_{i}&l\neq j\\[7.22743pt] \sum_{m=0}^{1}\gamma^{(i)}_{j,m}\,|\underline{j+m}\rangle\,|m\rangle_{i}&l=j\end{cases}\,;$ (A.2)
moreover
$I^{(i)}_{l}\,|\underline{j}\rangle\,|1\rangle_{i}=|\underline{j}\rangle\,|1\rangle_{i}\,,\qquad\qquad l\neq j-1\,.$ (A.3)
It follows that the $I^{(i)}_{l}$ operators do not interfere with each other:
@@
since $\mathcal{U}\,|\underline{0}\rangle\,|0\rangle^{\otimes n}=|\underline{k}\rangle\,|D^{n}_{k}\rangle$, see (3.13). It is possible to show that the product over $l$ in (A.1) can be restricted, so that $U_{i}$ is given by
@@
Let us represent $U_{i}$ (A.8) by the circuit diagram in Fig. 2(a), where the top wire represents qubit $i$, and the bottom wire represents the $(k+1)$-level qudit ancilla. The operator $\mathcal{U}$ (A.7) can then be represented by the circuit diagram in Fig. 2(b). The size and depth of the circuit is $3k(n+1-k)$, as follows from (A.7) and (A.8). Code in cirq [39] for simulating this circuit is provided in the Supplemental Material [40].
@@
[…]
Figure 2: Circuit diagram for sequential Dicke state preparation. (a) $U_{i}=\prod_{l}I^{(i)}_{l}$, with $x=\max(0,i-n+k-1)$ and $y=\min(i-1,k-1)$; (b) $\mathcal{U}=\prod_{i}U_{i}$